p(x) = q x, x = 1, 2, 3, 4.
Value of q
Probability distribution table of X
Mean, E(x)
E(x2)
Variance, σ2
Standard deviation, σ
1
Expert's answer
2019-08-19T10:23:29-0400
"\sum p(x_i)=1;
\\q(x_1+x_2+x_3+x_4)=1;
\\q(1+2+3+4)=1;
\\q=0.1;"

Using the formula p(x) and the resulting value q form a table of probability distribution


"\begin{matrix}
 x & 1 & 2 & 3 & 4 \\
 p(x) & 0.1 & 0.2 & 0.3 & 0.4;
\end{matrix}""E(x)=\sum p(x_i)x_i;
\\E(x)=0.1*1+0.2*2+0.3*3+0.4*4=3;"

"E(x^2)=\sum p(x_i)x_i^2
\\E(x^2)=0.1*1+0.2*4+0.3*9+0.4*16=10;"

"\sigma^2=E(x^2)-E(x)^2;
\\\sigma^2=10-9=1;"

"\sigma=\sqrt{E(x^2)-E(x)^2};
\\\sigma=\sqrt{10-9}=1;"